OpenStudy (anonymous):

how do you establish this identity (tan v+1)^2 +(tan v-1)^2 =2sec^2 v

OpenStudy (loser66):

expand the square, what do you get?

OpenStudy (anonymous):

(tan v^2 +2tan v +1)+(tan^2-2tan+1)= 2tan^2 +1

OpenStudy (anonymous):

+2 sorry

OpenStudy (loser66):

ok, factor 2 out, what do you get?

OpenStudy (anonymous):

tan^2+1=sec^2

OpenStudy (loser66):

2 out , so it is 2 sec^2, right? there you go,

OpenStudy (anonymous):

ok Thanks so much for your help :)

OpenStudy (loser66):

wut wut. glad to hear!!